TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes


Created by Main.BorisBellesia on 2007-01-31 - 20:38

SCT UA 67 - UJ 67, Jan. 31st.

PDF RAT_20070131.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x31x20x37/RAT_20070131.pdf

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-01-30 - 14:19

RAT in Point 8 - Tuesday 30/01

PDF RAT_20070130.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x30x14x18/RAT_20070130.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-30 - 12:35

SCT_UA 67- UJ 67, Jan. 30th 2006

PDF RAT_20070130.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x30x12x14/RAT_20070130.pdf

Comments:

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-29 - 17:06

RAT in Point 8 - Monday Jan 29

PDF RAT_20070129_withAttachments.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x29x17x02/RAT_20070129_withAttachments.pdf

  • The short to ground in Sector 78 disappeared last Saturday
  • The cooling capacity was restored to 80% today
  • The mobile camera run will be planned tomorrow from Point 7
  • Preparation for Powering Tests will take place from tomorrow in UA83

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-29 - 12:45

SCT_UA 67- UJ 67, Jan. 29th 2006

PDF RAT_20070129.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x29x12x44/RAT_20070129.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-01-26 - 18:03

Beam Dumping System_RAT - January, 26

PDF RAT_20070126.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x26x18x01/RAT_20070126.pdf

PDF LSS6_v1.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x26x18x01/LSS6_v1.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-26 - 16:43

SCT_UA 67- UJ 67, Jan. 26th 2006

PDF RAT_20070126.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x26x16x41/RAT_20070126.pdf

Comments:

  • At 15:00 8hr run of the 13 kA circuits have been STOPPED due to water cooling circuit.
  • TS-CV intervention scheduled for Monday 29.01
  • 8hr run for the 13 kA circuits foreseen for the next Tuesday (30.01) after TS-CV intervention.

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-26 - 12:27

RAT in point 8 - 26th January 2007

PDF RAT_20070126.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x26x12x26/RAT_20070126.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-01-25 - 18:49

Preparation for the Powering Tests Sector 7-8

PDF RAT_78_070124.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x25x18x48/RAT_78_070124.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-25 - 17:05

AUG scheduling meeting

PowerPoint Tests_annuels_des_AUG_LHC.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x25x16x36/Tests_annuels_des_AUG_LHC.ppt

Present :

P. Anderssen, B. Bellesia, C. Castillo Trello, G. Cumer, C. Dehavay, P. Fessia, J.P. Malod Dognin, M. Pojer, R. Saban, E. Barbero Soto, J.Toullieux, S. Weisz


* Presentation of G. Cumer (see attached file) - The procedures adopted during an AUG have been mentioned. - Resume of the affected Surface and Underground Areas during the AUG in Point 6.

* Friday 9 Feb. can be a possible date for a AUG in point 6: Waiting for the “ok” of Installation (S. Weisz/ TS-IC) and Interconnections (P. Fessia/ AT-MCS)

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-25 - 16:31

SCT_UA 67- UJ 67, Jan. 25th 2006

PDF RAT_20070125.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x25x16x30/RAT_20070125.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-01-25 - 11:47

Preparation for Cool-Down Sectors 7-8,4-5,8-1

PDF RAT_20070125_2.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x25x11x45/RAT_20070125_2.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-24 - 19:21

SCT_UA 67- UJ 67, Jan. 24th 2006

PDF RAT_20070124.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x24x19x21/RAT_20070124.pdf

Comments:

Edit, attach file or add comment


Created by Main.JeanPierreMalodDognin on 2007-01-24 - 16:56

la mise en froid se deroule tres normalement...

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-23 - 14:31

RAT in point 8 - 23th January 2007

PDF RAT_20070123.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x23x14x25/RAT_20070123.pdf

Comments:

  • NEXT RAT will be on thursday 25th january 2007 at 8:30 -- Main.BorisBellesia - 23 Jan 2007 - 14:31

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-01-22 - 17:17

RAT in Point 8 - 22nd January 2007

PDF RAT_20070122.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x22x17x13/RAT_20070122.pdf

Comments:

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-19 - 19:37

RAT in Point 8 - Minutes - 19th Jan 2007

PDF RAT_20070119.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x19x19x35/RAT_20070119.pdf

Edit, attach file or add comment


Created by Main.JeanPierreMalodDognin on 2007-01-19 - 16:45

Choses vues avant les tests...

JPEG IMGP9979.JPG:
IMGP9979.JPG

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-01-19 - 12:24

RAT for SCT in UA67/UJ67 - Jan, 19

PDF RAT_20070119.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x19x11x25/RAT_20070119.pdf

PDF SCT-Planning-UA67UJ67-v6.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x19x11x25/SCT-Planning-UA67UJ67-v6.pdf

Open Issues

  • 17.01 Broken thermostat on 13 kA EE
  • 10.01 QPS racks connection
  • 08.01 EE fan rotation problem – all ring
  • 21.12 Aluminum panel for the cables ducts/ shipping and install.

Heat Runs and Test Schedule

08.01 DONE 600A and 13 kA EE system IST.
08.01 DONE RB IST. ON GOING
11.01 DONE 120A powering tests.
15.01 DONE 2 kA powering tests.
15.01 DONE 600A powering tests.
17.01 DONE 600A-EE powering tests.
22.01 - PIC-PC-QPS Interface tests.
23.01 - RQD&F powering tests.
24.01 - RB powering tests.
26.01 - 120A-600A-2kA-RQD/F-RB 8h heat run.
05.02 - 24h heat run test.
07.02 - Polarity test.
14.02 - AUG&UPS test.
- Main.MirkoPojer - 19 Jan 2007 - 11:28

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-18 - 11:14

Cool-Down follow up - Jan 18th - Plots

PDF ELQA-HVQ_cooldown_2007-01-18.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x18x10x46/ELQA-HVQ_cooldown_2007-01-18.pdf

PDF Cooling_LHC_sector78_18Jan07.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x18x10x46/Cooling_LHC_sector78_18Jan07.pdf

Comments:

  • Inlet set temperature = 185K
  • Cooling Power = 300kW = 50% of nominal
  • Leakage Current Monitoring = ok

MircoCoccoli - 18 Jan 2007 - 10:51

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-17 - 17:24

RAT in point 8 - 17th January - Minutes

PDF RAT_20070117_with_NC_Mag_Planning.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x17x17x17/RAT_20070117_with_NC_Mag_Planning.pdf

> ATTENTION: next RAT in point 8 on FRIDAY 19th at 8.30am <

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-17 - 16:43

Minutes of the meeting on Point 8 Warm Magnets Commissioning - 15th Jan

PDF Minutes_WarmMag_LSS8_15Jan07.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x17x15x56/Minutes_WarmMag_LSS8_15Jan07.pdf

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-17 - 11:24

SCT_UA 67- UJ 67, Jan. 17th 2007

PDF RAT_20070117.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x17x11x19/RAT_20070117.pdf

Open Issues

  • 17.01 Broken thermostat on 13 kA EE
  • 12.01 Power cut in ERD-R13
  • 12.01 Cleaning of filters by CV (planned for 17.01)
  • 10.01 QPS racks connection
  • 08.01 EE fan rotation problem – all ring
  • 21.12 Aluminum panel for the cables ducts/ shipping and install.
  • 21.12 HV Insulation test of the DQS poles

Heat Runs and Test Schedule

08.01 ON GOING 600A and 13 kA EE system IST
08.01 ON GOING RB IST
11.01 DONE 120A powering tests.
15.01 DONE 2 kA powering tests.
17.01 ON GOING PIC-PC-QPS Interface tests.
19.01 DONE 600A powering tests.
23.01 - 600A-EE powering tests.
26.01 - 120A-600A-2kA 8h heat run.
30.01 - RQD&F powering tests.
02.02 - RB powering tests.
07.02 - RQ&RB 8h heat run.
12.02 - 24h heat run test.
14.02 - AUG&UPS test.
14.02 - Polarity test.

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-16 - 18:33

RAT in point 8 - 16th January - Minutes

PDF RAT_20070116.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x16x18x25/RAT_20070116.pdf

Open Issues:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Instrumentation Cabling
  • 12.01.07 Cryo-instrumentation cables: 2 LSS8L connectors (A.Suraci-J-C.Guillaume)
  • 10.01.07 Continuity error with cable 1813004A and a wrong housing of 1702440 which should be replaced (D.Bozzini)
Cryogenics
  • 16.01.07 Vac-Cryo Interlock tests – window of 1 day needed (P.Cruikshank-F.Millet)
  • 22.11.06 Instrumentation and remote control not fully stable (P.Gayet - F.Millet)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 08.01.07 Re-installation of 3 BLMs interfering with pumping groups (B.Dehning)
  • 13.11.06 BPM connection in Q2 (R.Jones) – waiting for Inner Triplet to be repaired

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-01-16 - 10:17

First Cooldown QRL C to D

Images obtained from Francois Millet

PNG Cooldown.png:
Cooldown.png
First cooldown QRL 78
PNG First_cooldown_Q9R7.png:
First_cooldown_Q9R7.png
First cooldown Q9R7

Comments:

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-15 - 18:42

RAT Meeting in Point 8 - Jan 15th

PDF RAT_20070115.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x15x18x38/RAT_20070115.pdf

MircoCoccoli - 15 Jan 2007 - 18:40

Open Issues:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Instrumentation Cabling
  • 12.01.07 Cryo-instrumentation cables: 2 LSS8L connectors (A.Suraci-J-C.Guillaume)
  • 10.01.07 Continuity error with cable 1813004A and a wrong housing of 1702440 which should be replaced (D.Bozzini)
Cryogenics
  • 22.11.06 Instrumentation and remote control not fully stable (P.Gayet - F.Millet)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 08.01.07 Re-installation of 3 BLMs interfering with pumping groups (B.Dehning)
  • 13.11.06 BPM connection in Q2 (R.Jones) – waiting for Inner Triplet to be repaired

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-01-15 - 11:21

RAT for SCT in UA67/UJ67 - Jan, 15

PDF RAT_20070115.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x15x11x07/RAT_20070115.pdf

Open Issues:

  • 15.01 Loosen bolts on RCBYH4.R6B1 and RCBYV4.R6B2 PCs.
  • 12.01 Power cut in ERD-R13
  • 12.01 Cleaning of filters by CV (planned for 17.01)
  • 10.01 QPS racks connection
  • 21.12 Aluminum panel for the cables ducts/ shipping and install.
  • 21.12 HV Insulation test of the DQS poles
Heat Runs and Test Schedule:
  • 08.01 600A and 13 kA EE system IST. ON GOING
  • 08.01 RB IST. ON GOING
  • 11.01 120A powering tests. COMPLETED
  • 15.01 2 kA powering tests. ON GOING
  • 17.01 PIC-PC-QPS Interface tests.
  • 19.01 600A powering tests. COMPLETED
  • 23.01 600A-EE powering tests.
  • 26.01 120A-600A-2kA 8h heat run.
  • 30.01 RQD&F powering tests.
  • 02.02 RB powering tests.
  • 07.02 RQ&RB 8h heat run.
  • 12.02 24h heat run test.
  • 14.02 AUG&UPS test.
  • 14.02 Polarity test.

- Main.MirkoPojer - 15 Jan 2007 - 11:15

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-12 - 18:46

RAT Meeting in Point 8 - Jan 12th

PDF RAT_20070112.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x12x18x43/RAT_20070112.pdf

MircoCoccoli - 12 Jan 2007 - 18:45

Open Issues (Sect 78):

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL & VAC.
DC cabling
  • 10.01.07 Connection of the NC Magnets in LSS8R (J-C.Guillaume)
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL) – PC Non conformities (H.Thiesen)
Instrumentation Cabling
  • 12.01.07 Cryo-instrumentation cables: 2 LSS8L connectors (A.Suraci-J-C.Guillaume)
  • 10.01.07 Continuity error with cable 1813004A and a wrong housing of 1702440 which should be replaced (D.Bozzini)
Cryogenics
  • 22.11.06 Instrumentation and remote control not fully stable (P.Gayet - F.Millet)
DFB Commissioning
  • 29.11.06 DFBX Pumping of Current Leads insulation vacuum (S.Feher)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1 (D.Bozzini)
AOB
  • 08.01.07 Re-installation of 3 BLMs interfering with pumping groups (B.Dehning)
  • 13.11.06 BPM connection in Q2 (R.Jones) – waiting for Inner Triplet to be repaired

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-01-12 - 15:48

RAT for SCT in UA67/UJ67 - Jan, 12

PDF RAT_20070112.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x12x15x43/RAT_20070112.pdf

Open Issues:

  • 12.01 Power cut in ERD-R13
  • 12.01 Cleaning of filters by CV (planned for 17.01)
  • 10.01 QPS racks connection
  • 08.01 EE fan rotation problem to be fixed for the whole machine
  • 21.12 Aluminum panel for the cables ducts/ shipping and install.
  • 21.12 HV Insulation test of the DQS poles
Heat Runs and Test Schedule:
  • 08.01 600A and 13 kA EE system IST. ON GOING
  • 08.01 RB IST.
  • 11.01 120A powering tests.
  • 15.01 2 kA powering tests.
  • 17.01 PIC-PC-QPS Interface tests.
  • 19.01 600A powering tests.
  • 23.01 600A-EE powering tests.
  • 26.01 120A-600A-2kA 8h heat run.
  • 30.01 RQD&F powering tests.
  • 02.02 RB powering tests.
  • 07.02 RQ&RB 8h heat run.
  • 12.02 24h heat run test.
  • 14.02 AUG&UPS test.
  • 14.02 Polarity test.

-- Main.MirkoPojer - 12 Jan 2007 - 15:46

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2007-01-11 - 18:27

RAT Meeting P8 - 11th Jan Minutes

PDF RAT_20070111.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x11x18x26/RAT_20070111.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-10 - 17:47

SCT_UA 67- UJ 67, Jan. 10th 2007

PDF RAT_20070110.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x10x17x43/RAT_20070110.pdf

Comments:

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2007-01-10 - 15:13

RAT Meeting P8 - 10th Jan Minutes

PDF RAT_20070110.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x10x15x12/RAT_20070110.pdf

Comments:

Edit, attach file or add comment


Created by Main.IvanMoyaMartinez on 2007-01-09 - 11:05

Labels of the Water cooled cables following the ECR- EDMS 804188

Some pictures of the water-cooled cables with the labels showing they are following the ECR- EDMS 804188, when it's necessary

JPEG P1000209.JPG:
P1000209.JPG

JPEG P1000206.JPG:
P1000206.JPG

JPEG P1000207.JPG:
P1000207.JPG

JPEG P1000208.JPG:
P1000208.JPG

Comments:

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2007-01-08 - 20:35

RAT Point 8 Minutes

PDF RAT_20070108.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x01x08x20x33/RAT_20070108.pdf

RAT regarding Cool Down of Sect 78 and preparation for Cool Down of sectors 45 and 81

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-01-08 - 18:46

SCT_UA 67- UJ 67, Jan. 8th 2007

PDF RAT_20070108.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x01x08x18x45/RAT_20070108.pdf

Comments:

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-01-08 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login